OpenStudy (code45):

dB=10log(P2/5) Solve for P2.

OpenStudy (anonymous):

not sure if i got it correct but my solution is this: dB = 10log(P2/5) dB/10 = log(P2/5) 10^(dB/10) = P2/5 (5)(10^(dB/10))=P2